RAL Tier1 Experiments Liaison Meeting
Access Grid
RAL R89
-
-
13:30
→
13:31
Experiment Operational Issues 1m
-
13:35
→
13:40
ATLAS Operations Report 5mSpeakers: Brij Kishor Jashal (Rutherford Appelton Laboratory), Jyoti Prakash Biswal (Rutherford Appleton Laboratory)
-
13:40
→
13:45
CMS Operations Report 5mSpeaker: Katy Ellis (Science and Technology Facilities Council STFC (GB))
Spotted 25% of CMS transfer from Echo to Antares failing. ATLAS had the same. Only IPv4 transfers were failing and not all of them. IPv6 was 100% successful. At first guessed it was related to Antares upgrade/new hardware, but then asked Jyothish and he quickly proved it was missing IPv4 routes to Antares from the new 23 tranche of Echo gateways. These are now blacklisted and Jyothish is following up with DI. In the meantime the failures stopped.
There is clearly Antares data moving using IPv6. But because the current MGM does not have a publicly routable IPv6 address (quote: George) the CMS SAM tests remain in warning state.
There have been some issues with the new ETF tests (token) for ARC-CEs, which were briefly included in the Site Status. These have now been removed from Site Status until better stability is achieved. Errors have been 'corrected' by CMS.
The possible network issue stopping jobs being submitted from Schedds at FNAL seems to have been solved. So far as we know nothing was changed on the RAL or the FNAL side. But we can now do traceroute etc. to the relevant machine(s).
-
13:45
→
13:50
LHCb Operations Report 5mSpeaker: Alexander Rogovskiy (Rutherford Appleton Laboratory)
-
13:50
→
13:55
ALICE Operations Report 5mSpeaker: Alexander Rogovskiy (Rutherford Appleton Laboratory)
-
13:55
→
14:00
LSST Operations Report 5mSpeakers: Mathew Sims, Timothy Noble (Science and Technology Facilities Council STFC (GB))
- ComCam data (ccms1) now at RAL as well as being moved from RAL to Lancs as well
- Being ingested into the metadata service (Butler now)
- My anticipation is that this dataset will be run in August after the LSST Community Workshop next week
- RC2 and DC2 data retrieved to US and confirm configuration and return processes
- Now to run a coordinated DC2 and RC2 run at all 4 sites (4x2) return all the data and investigate results for deviation
- Will move message stack from Cloud to VMWare / Proxmox to ensure stability of this going forwards
- Working with Edinburgh to facilitate the set up of the access centre for astronomers
- ComCam data (ccms1) now at RAL as well as being moved from RAL to Lancs as well
-
14:00
→
14:01
Tier-1 Projects 1m
-
14:15
→
14:25
Anatares Upgrade 10m
New EOS nodes
Repack ProgressSpeakers: George Patargias, Thomas Byrne - 14:25 → 14:35
-
14:35
→
14:45
Utilizing GPUs 10mSpeakers: Jyoti Prakash Biswal (Rutherford Appleton Laboratory), Thomas Birkett
-
14:45
→
14:50
SSD Storage Evaluation 5m
-
14:50
→
14:55
Echo deployment 5m
-
15:00
→
15:01
AOB 1m
-
15:01
→
15:10
Summary of Operational Status and Issues 9mSpeakers: Brian Davies (Lancaster University (GB)), Darren Moore
-
15:10
→
15:15
Any other Business 5mSpeakers: Brian Davies (Lancaster University (GB)), Darren Moore
-
13:30
→
13:31